OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. (2 points)

OpenStudy (anonymous):

y = mx + b m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

@lolly wanna try

OpenStudy (anonymous):

no..but i will..:)

OpenStudy (anonymous):

(1,-4), m = -4 y = mx + b Sub to find b -4 = -4*1 + b b = 0 ------- y = -4x m = -4 means slope = -4. slope-intercept form of equation of a straight line is y = mx + b. m is the slope. b is the y-intercept. your equation starts of with the slope-intercept generic form of y = mx + b replace m with -4 to get y = -4x + b your line goes through the point (x,y) = (1,-4) that means the x-coordinate is 1 and the y coordinate is -4. to find the y-intercept, you replace y with -4 and x with 1 in the equation and solve for b. your equation of y = -4x + b becomes: -4 = (-4 * 1) + b

OpenStudy (anonymous):

can you set up the equation? then ill solve:P

OpenStudy (anonymous):

@swiftskier96 is that the answer http://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.233716.html hahaha

OpenStudy (anonymous):

(x1,y1) = (–1, –4) (x2,y2) = (1, 4)

OpenStudy (anonymous):

soo y=mx=b? isnt that for finding slope?

OpenStudy (anonymous):

yes but you need to find m 1st \[y=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

meant m =

OpenStudy (anonymous):

ook ill try it..hold on..:)

OpenStudy (anonymous):

m=-6

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

show me how?

OpenStudy (anonymous):

m is wrong :D

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

m= 1-4/-1--4. m=-3/3 m=-6

OpenStudy (anonymous):

(x1,y1) = (–1, –4) (x2,y2) = (1, 4) x1 = -1 y1 = -4 x2 = 1 y2 = 4

OpenStudy (anonymous):

hi lolly :P

OpenStudy (anonymous):

right..thats what i did..,.i think..?

OpenStudy (anonymous):

hi..:)

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1)

OpenStudy (anonymous):

@surdawi how did i do it rong?

OpenStudy (anonymous):

(4- - 4)/(1- -1)= 8/2=4

OpenStudy (anonymous):

so m is 4?

OpenStudy (anonymous):

what you did (x2-y2)/(x1-y1)

OpenStudy (anonymous):

well i thought it was -6..but its not..?

OpenStudy (anonymous):

m = 4 now you choose wither point and solve for b y = 4x + b

OpenStudy (anonymous):

k hold on ill try again..:)

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }=\frac{ 4- -4 }{ 1- -1 }=\frac{ 8 }{ 2 }=\]

OpenStudy (anonymous):

for slope slope=(y2-y1)/(x2-x1) it can be directly solve see the formula that i have given already

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1)

OpenStudy (anonymous):

oooh!!! i got the x1and 2s switched up..ok i see..:)

OpenStudy (anonymous):



OpenStudy (anonymous):

the one with less numbers right?

OpenStudy (anonymous):

there is no zeros for x or y this time so pick one and solve, it doesnt matter which one

OpenStudy (anonymous):

which one would be easier?

OpenStudy (anonymous):

either, same thing

OpenStudy (anonymous):

k ill do the one with less numbers..:P

OpenStudy (anonymous):

so 4=4x1=b? 4=b..but we already knew that..

OpenStudy (anonymous):

both of them has same amount of numbers :p 4=4x1=b? ! why is there an equal to b

OpenStudy (anonymous):

oooh! whoop..my bad..:)

OpenStudy (anonymous):

so 4=4+4...

OpenStudy (anonymous):

4=8? that dosent make sense...

OpenStudy (anonymous):

4=4x1+ b what is x1 = ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

y1 =?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

000h!!!

OpenStudy (anonymous):

no it dosent..:)

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

it should be one of the pairs either (1,4) or (-1,-4) -4 = 4*(-1) + b

OpenStudy (anonymous):

-4=4x-1=4?

OpenStudy (anonymous):

4*-1=?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

-4 = -4 + b how can you keep b on one side?

OpenStudy (anonymous):

by adding 4 to b?

OpenStudy (anonymous):

ok show me how?

OpenStudy (anonymous):

-4=-4+b (then the 4 crosses out) +4 +4

OpenStudy (anonymous):

-4 +4 = -4 + 4 +b 0 = b y = 4x + 0

OpenStudy (anonymous):

where did the other set of -4=4 come from?

OpenStudy (anonymous):

you have to add 4 to both sides

OpenStudy (anonymous):

passes through points (–1, –4) and (1, 4)?

OpenStudy (anonymous):

Anamaria what grade you in?

OpenStudy (anonymous):

10.:)

OpenStudy (anonymous):

you should know how to add :P

OpenStudy (anonymous):

lol i do know how to add:)

OpenStudy (anonymous):

-4=-4+b (then the 4 crosses out) +4 +4 was this rong?

OpenStudy (anonymous):

lol @surdawi u sleeping?

OpenStudy (anonymous):

i went to pray